calculate the ph 0f final solution of following
20 ml of 0.1M ch3cooh,Ka 1.8*10^-5

Grade:12th pass

1 Answers

Vikas TU
6 years ago
the a Ph becomes:
                     CH3COOH (aq) + H2O (l)   CH3COO-(aq)+ H3O+(aq)
 Ions        0.1M
                -G                               +G                  +G
                          0.1 M    -G                               +G                  +G 
dissociation constant : Ka= 1.8*10^-5
                                                                        [CH3COOH][H3O+][CH3COO]=1.8x10^5[CH3COOH]
Replacing with G
         We have [G][G]/0.1-G =1.8x10^-5
Solving using the quadratic formula : G = 3.399 x 10^-5
                                                        pH = -log(G) = 4.4686